OpenStudy (clarence):

The line x+7y = 20 is a normal to the curve y = x^3+ax^2+bx+5 at the point where x = -1, when determining the slope of the tangent to this point, which equation am i suppose to use??

OpenStudy (anonymous):

FInd the gradient of the line x+7y = 20.

OpenStudy (clarence):

so, 7y=21, and y =3? :)

OpenStudy (anonymous):

wht is the question?

OpenStudy (clarence):

The line x +7y =20 is a normal to the curve y = x^3+ax^2+bx+5 at the point where x = -1. Determinet the slope of the tangent at this point.

OpenStudy (anonymous):

do you know this one slope of tangent * slope of normal = -1

OpenStudy (clarence):

you can teach me?

OpenStudy (anonymous):

okay :) this is a rule such that at a point in a curve, the gradient/slope of tangent at a point times the gradient/slope of a normal at that point is equal to -1

OpenStudy (anonymous):